# Chebyshev polynomials and a refinement of the local residue/non-residue structure at a prime

### v1

@@ Abstract.
The basic power function $t_{n}(x)=x^{n}$ is in some sense a classical limit for large $x$, of the monictised Chebyshev polynomial of the first kind $T_{n}(x)/2^{n-1}$. A theorem of Ritt says they are the only two families of polynomials $p_{n}(x)$ over $\mathbb{C}$ which satisfies the commutativity relation $p_{n}(p_{m}(x))=p_{m}(p_{n}(x))$. The commutativity $t_{n}(t_{m}(x))=t_{m}(t_{n}(x))$ is the reason why the RSA scheme allow also digital signature but the Diffie-Hellman key exchange protocol depends only on the commutativity. The DH scheme and many results in elementary local (at a fixed prime) multiplicative number theory is about properties of the power function $t_{n}(x)$ and they have natural analogue extension to $T_{n}(x)$. Recently we discovered a Chebyshev version of Euler’s primality criterion , which however depends on two quadratic characters $\epsilon_{p}(a)=\left(\frac{a^{2}-1}{p}\right)$ and $\delta_{p}(a)=\left(\frac{2(a+1)}{p}\right)$. This gives rise to a local partition of the integers mod an odd prime $p$ in two distinct ways into 4 disjoint sets $A_{\epsilon\delta}$ of smaller size about $p/4$ over which the exponential sum still have squar root cancellations. This can be thought of as a real refinement of the residue/non-residue as it arise from viewing $T_{n}(x)$ is the ”real” part of the $n$th power of the unit $\omega_{x}=x+\sqrt{x^{2}-1}$, namely $\omega_{x}^{n}=T_{n}(x)+U_{n-1}(x)\sqrt{x^{2}-1}$. There are obvious analogue of Chebyshev version of pseudoprimes, Wieferich primes, Lucas-Lehmer, AKS, Diffie-Hellman, cyclotomic expansions and probably many others which hopefully some may give improved resuts on the original version.
@@ 1. Introduction
at any odd prime $p$. In fact we have
@@ 2. A Chebyshev-Euler primality criteria
Here the character $\epsilon(a)$ comes from our qaryq-ary extension of Lucas -Lehmer [3], based on numerical observation. The second character $\delta$ comes out when one follows the proof of the $2$-ary Lucas-Lehmer necessity condition for primality of Mersenns primes [6]. One just replace $\omega_{2}=2+\sqrt{3}$ by $\omega_{a}$ and carry along the argument with the character $\epsilon$.
@@ 2.1. Chebyshev-Wieferich primes
This is equivalent to $T_{(p-\epsilon)/2}(a)=\delta\mod p^{2}$ and $U_{(p-\epsilon)/2-1}(a)=0\mod p^{2}$, but by (2.4), we know the first condition already holds for any prime $p$ so that $p$ is a Chebyshev-Wieferich prime if and only if $U_{(p-\epsilon)/2-1}(a)=0\mod p^{2}$. Also a Chebyshev-Wieferich prime is automatically a CW-plus prime, namely $U_{(p-\epsilon)/2}(a)=0\mod p^{2}$ implies
@@
I have seen the prime 103, 1039 before (especially 1039 because it is not quite 1093 as some kind of Wieferich primes but I can’t recall where now.
@@ 3.4. Commutativity and Chebyshev-Diffie-Hellman key exchange protocol
Recall that a public key encrytion scheme use a pair of keys $(d,e)$. Every one can encrypt a message $x$ using the public key $e$ and and encryption algorithm $E_{e}(x)$ but only the holder of the secret key $d$ can decrypt by computing $D_{d}(E_{e}(x))=x$. The first successful implementation of such a PK system is the RSA [7] scheme which use the same algorithm for $D$ and $E$ ie. $E_{e}(x)=t_{e}(x)\mod m$, $D_{d}(x)=t_{d}(x)\mod m$ where key pairs satisfies $de=1\mod\phi(m)$. Because of the commutativity $t_{d}(t_{e}(x))=t_{e}(t_{d}(x))$ we have $E_{e}(D_{d}(x)=D_{d}(E_{e}(x))=x$. This means the holder of the secret key can ”sign” a $x$ now thought of as a document by decrypting it to produce a signature $s=D_{e}(x)$ which the public can verify its authenticity by encrypting to recover the original document $x=E_{e}(s)=E_{e}(D_{d}(x))$. Since only the holder of the secret key can produce a signature which can be encrypted/verified to recover a meaningful $x$, this prove he/she must be the real signer. So the commutativity is the reason why RSA enable also digital signature. However it seems not possible to construct a Chebyshev-RSA scheme since we also need a homomorphism property $t_{n+m}(x)=t_{n}(x)t_{m}(x)$ which is not available for Chebyshev. However the Diffie-Hellman key exchange protocal uses only the commutativity property and seemingly the Chebyshev discrete logarithm problem : finding $n$ which satisfies $T_{n}(a)=m\mod p$ for a large prime $p$, given $a,m$ is also hard.
@@ 3.5.1. Square root cancellation in half subset of the quadratic residue or non residue
$Q=[1,4,9,16,2,13,3,18,12,8,6]$
$Q+1=[2,5,10,17,3,14,4,19,13,9,7]$
The symmetric elements of $Q+1$ is thus
@@
$[9,12,13,8,3,18],[1,4,16,2,6]$
We also have
Q1=[0,3,8,15,1,12,2,17,11,7,5]Q-1=[0,3,8,15,1,12,2,17,11,7,5]
@@
$[1,9,16,12,13],[4,3,18,8,6].$
Note $A_{++}$ appears in both and $A_{+-}$ does not appear above. It will appear twice when we shift $N\pm 1$. It is always the case that $R\pm 1,N\pm 1$ is essentially a union of two of the $A_{\epsilon\delta}$ (drop the 1,p11,p-1). Since 0 is symmetric and $\epsilon(0)=\left(\frac{-1}{p}\right)$, the symmetric sets are $A_{+1},A_{++}$ with $\epsilon=+$ when $p=1\mod 4$ and $A_{--},A_{-+}$ when $p=3\mod 4$.
So indeed we have a refinement of the residue/nonresidue which preserve provable square-root cancelations. We may also verify (2.8) for $p=23$ above, we must have
@@
We can also verify that $ord_{p}(\omega_{a})=24$ for the unreal primitive root $a=19$ . This order is the least $n$ such that $\omega_{a}^{n}=1\mod p$ or the least $n$ such that $([T_{n}(a),U_{n-1}(a)]=[1,0]\mod p$. It also equals the period of the second order linear recurrence $T_{0}(a)=1,T_{1}(a)=a,T_{n+1}(a)=2aT_{n}(a)-T_{n-1}(a)\mod p$ which is the least $n$ such that $T_{n}(a)=1,T_{n-1}(a)=T_{n+1}(a)=a\mod p$ but it is given just by the least $n$ such that $T_{n}(a)=1\mod p$ as it is automatic that $T_{n}(a)=1\mod p$ implies $T_{n+1}(a)=a,U_{n-1}(a)=0\mod p$, so we need only compute the first coordinate below
@@
As noted before even though $19$ is a primitive root only half the subset of $R_{p}$ appears in the first coordinate, only those with $\epsilon(a)=-1,0$ and note the first coordinate is palidormic, as this is a negative cycle as $I_{24}\subset A_{--}$. Actually this is some kind of second order Lucas sequence for fixed $a=19$ which is well known. Our main observation is that there is a structure to view all the $a$ together, $R_{p}$ is like a moduli space for all the order $(\mathbb{Z}/p\mathbb{Z})[\sqrt{a^{2}-1}]$.
@@ 3.6. Chebyshev-AKS
$F(y):=\sum_{t\geq 0}{n\choose 2t}y^{t}=\frac{(1+sqrt{y})^{n}+(1-\sqrt{y})^{n}}{2}$
@@
$T_{n}(x)=2^{n-1}x^{n}+\sum_{k=1}^{\lfloor n/2\rfloor}(-1)^{k}{n-k-1\choose k-1}\frac{n}{k}2^{n-(2k+1)}x^{n-2k}.$
If we let $A_{n,k}=\frac{a_{k}}{n}=(-1)^{k}{n-k-1\choose k-1}\frac{1}{k}2^{n-(2k+1)}$ be the coefficient of the RHS divided by $n$. Then clear $A_{n,n/2}=\frac{1}{n}$ for even $n$. If $n$ is odd composite and $p$ is any prime dividing $n$, then since $p$ does not divide ${n-p-1\choose p-1}=1\mod p$, $a_{p}\not=0\mod n$. ∎

### v2

@@ Abstract.
The basic power function $t_{n}(x)=x^{n}$ is in some sense a classical limit for large $x$, of the monictised Chebyshev polynomial of the first kind $T_{n}(x)/2^{n-1}$. A theorem of Ritt says they are the only two families of polynomials $p_{n}(x)$ over $\mathbb{C}$ which satisfies the commutativity relation $p_{n}(p_{m}(x))=p_{m}(p_{n}(x))$. The commutativity $t_{n}(t_{m}(x))=t_{m}(t_{n}(x))$ is the reason why the RSA scheme allow also digital signature but the Diffie-Hellman key exchange protocol depends only on the commutativity. The DH scheme and many results in elementary local (at a fixed prime) multiplicative number theory is about properties of the power function $t_{n}(x)$ and they have natural analogue extension to $T_{n}(x)$. Recently we discovered a Chebyshev version of Euler’s primality criterion , which however depends on two quadratic characters $\epsilon_{p}(a)=\left(\frac{a^{2}-1}{p}\right)$ and $\delta_{p}(a)=\left(\frac{2(a+1)}{p}\right)$. This gives rise to a local partition of $(\mathbb{Z}/p\mathbb{Z})\setminus\{\pm 1\}$ into 4 disjoint sets $A_{\epsilon\delta}$. This can be thought of as a real refinement of the residue/non-residue as it arise from viewing $T_{n}(x)$ is the ”real” part of the $n$th power of the unit $\omega_{x}=x+\sqrt{x^{2}-1}$, namely $\omega_{x}^{n}=T_{n}(x)+U_{n-1}(x)\sqrt{x^{2}-1}$. There are obvious analogue of Chebyshev version of pseudoprimes, Wieferich primes, Lucas-Lehmer, AKS, Diffie-Hellman, cyclotomic expansions and probably others.
@@ 1. Introduction
at any odd prime $p$. In fact we have for $n>2$,
@@ 2. A Chebyshev-Euler primality criteria
Here the character $\epsilon(a)$ comes from our qaryq-ary extension of Lucas -Lehmer [3], based on numerical observation. The second character $\delta(a)$ comes out when one follows the proof of the $2$-ary Lucas-Lehmer necessity condition for primality of Mersenns primes [7]. One just replace $\omega_{2}=2+\sqrt{3}$ by $\omega_{a}$ and follow the argument carring the character $\epsilon(a)$ along, and the character $\delta(a)$ appears in the final step just after (2.12) in the proof below.
@@ 2.1. Chebyshev-Wieferich primes
This is equivalent to $T_{(p-\epsilon)/2}(a)=\delta\mod p^{2}$ and $U_{(p-\epsilon)/2-1}(a)=0\mod p^{2}$, but by (2.4), we know the first condition already holds for any prime $p$ so that $p$ is a Chebyshev-Wieferich prime if and only if $U_{(p-\epsilon)/2-1}(a)=0\mod p^{2}$. Also a Chebyshev-Wieferich prime is automatically a CW-plus prime, namely $U_{(p-\epsilon)/2-1}(a)=0\mod p^{2}$ implies
@@
It seems we have seen the prime 103, 1039 before (especially 1039 because it is not quite 1093 as Wieferich primes of probably Lucas
@@ 3.3. Chebyshev cyclotomic expansion
There ia also a refined version of the cyclotomic expansion, since
$A_{\epsilon\pm}=\left\{a\in R_{p}:T_{(p-\epsilon)/2}(a)=\pm 1\right\},$
(2.8) above is equivalent to for $\epsilon=\pm 1$
$T_{(p-\epsilon)/2}(x)-1=(x-1)(2x+2)^{\lambda}\prod_{\begin{subarray}{c}d>2\\ d|(p-\epsilon)/2\end{subarray}}\Phi^{+}_{d}(2x)^{2}$
[…]
@@ 3.4. Commutativity and Chebyshev-Diffie-Hellman key exchange protocol
Recall that a public key encryption scheme [5] use a matching pair of keys $(d,e)$. Every one can encrypt a message $x$ using the public key $e$ and an encryption algorithm $E_{e}(x)$ but only the holder of the secret key $d$ can decrypt by computing $D_{d}(E_{e}(x))=x$. The first and still the most successful implementation of such a PK system is the RSA [8] scheme which use the same algorithm for $D$ and $E$ ie. $E_{e}(x)=t_{e}(x)\mod m$, $D_{d}(x)=t_{d}(x)\mod m$ where key pairs satisfies $de=1\mod\phi(m)$
Because of the commutativity $t_{d}(t_{e}(x))=t_{e}(t_{d}(x))$ we have
$E_{e}(D_{d}(x)=D_{d}(E_{e}(x))=x.$
This means the holder of the secret key can ”sign” a $x$ now thought of as a document by decrypting it to produce a signature $s=D_{e}(x)$ which the public can verify its authenticity by encrypting to recover the original document $x=E_{e}(s)=E_{e}(D_{d}(x))$. Since only the holder of the secret key can produce a signature which can be encrypted/verified to recover a meaningful $x$, this prove he/she must be the real signer. So the commutativity is the reason why RSA enable also digital signature. However it seems not possible to construct a Chebyshev-RSA scheme since we also need a homomorphism property $t_{n+m}(x)=t_{n}(x)t_{m}(x)$ which is not available for Chebyshev. Even though the above is well-known material, we cannot resist writing about it to point out the fact that commutativity is the reason that made digital signature possible and this seems to be still the most beautiful application of number theory.
However the Diffie-Hellman key exchange protocal uses only the commutativity property and seemingly the Chebyshev discrete logarithm problem : finding $n$ which satisfies $T_{n}(a)=m\mod p$ for a large prime $p$, given $a,m$ is also hard.
@@ 3.5.1. Square root cancellation in half subset of the quadratic residue or non residue
$R=[1,4,9,16,2,13,3,18,12,8,6]$
$R+1=[2,5,10,17,3,14,4,19,13,9,7]$
The symmetric elements of $R+1$ is thus
@@
$SR^{+}=[9,12,13,8,3,18],CR^{+}=[1,4,16,2,6],$
where here $SR^{+}$ means the symmetric subsets of $R$ partition via a +1+1 shift, and clearly reading off directly from $R$,
$SR^{+}=\{a\in R:p-2-a\in R\},CR^{+}=R\setminus SR^{+}$
and similarly using a negative shift,
$SR^{-}=\{a\in R:p+2-a\in R\},CR^{-}=R\setminus SR^{-},$
and similarly a partition of $N$, $SN^{\pm},CN^{\pm}$ using a $\pm 1$ shift of $N$. We also have
R1=[0,3,8,15,1,12,2,17,11,7,5]R-1=[0,3,8,15,1,12,2,17,11,7,5]
@@
$SR^{-}=[1,9,16,12,13],CR^{-}=[4,2,3,18,8,6].$
Note that we always have $1\in SR^{-}$ since $1\in R$ and $p+2-1=1\mod p$, and p1p-1 is always in $SR^{+}$ or $SN^{+}$ since $p-2-(p-1)=p-1\mod p$. Since the non-residues
$N=\{4,6,9,10,13,14,16,18,19,20,21\},$
we can just read off the symmetric partition
$SN^{+}=\{7,14,10,11,22\},SN^{-}=\{5,20,10,15,11,14\}.$
Note $A_{++}$ appears in both the $R\pm 1$ shift and $A_{+-}$ does not appear above. It will appear twice when we shift $N\pm 1$. It is always the case that $R\pm 1,N\pm 1$ is essentially a union of two of the $A_{\epsilon\delta}$ (drop the 1,p11,p-1). Since 0 is symmetric and $\epsilon(0)=\left(\frac{-1}{p}\right)$, the symmetric sets are $A_{+1},A_{++}$ with $\epsilon=+$ when $p=1\mod 4$ and $A_{--},A_{-+}$ when $p=3\mod 4$.
So indeed we have a refinement of the residue/nonresidue which preserve provable square-root cancelations.
We may also verify (2.8) for $p=23$ above to define the $A_{\epsilon\delta}$ via the multiplicative order, we must have
@@
We can also verify that $ord_{p}(\omega_{a})=24$ for the unreal primitive root $a=19$ . This order is the least $n$ such that $\omega_{a}^{n}=1\mod p$ or the least $n$ such that $([T_{n}(a),U_{n-1}(a)]=[1,0]\mod p$. It also equals the period of the second order linear recurrence $T_{0}(a)=1,T_{1}(a)=a,T_{n+1}(a)=2aT_{n}(a)-T_{n-1}(a)\mod p$ which is the least $n$ such that $T_{n}(a)=1,T_{n-1}(a)=T_{n+1}(a)=a\mod p$ but it is given just by the least $n$ such that $T_{n}(a)=1\mod p$ as it is automatic that $T_{n}(a)=1\mod p$ implies $T_{n+1}(a)=a,U_{n-1}(a)=0\mod p$, so we need only compute the first coordinate below. It seems interesting that the multiplicatively defined order can be computed additively as a period of a linear recurrence.
@@
As noted before even though $19$ is a primitive root only half the subset of $R_{p}$ appears in the first coordinate, only those with $\epsilon(a)=-1,0$ and note the first coordinate is palidormic, as this is a negative cycle as $I_{24}\subset A_{--}$. Actually this is some kind of second order Lucas sequence for fixed $a=19$ which is well known. Our main observation is that there is a structure to view all the $a$ together, $R_{p}$ is like a moduli space for all the order $(\mathbb{Z}/p\mathbb{Z})[\sqrt{a^{2}-1}]$. It seems useful to state our observation on the partition of the residue/non-residue.
Lemma 3.7.
Let $X$ be either $R$ or $N$. Then can can read off a partition of $X$ into two disjoint sets in two distinct ways which are translates of $A_{\epsilon\delta}$ shifted by $\pm 1$. WE define
$SX^{+}=\{a\in X:p-2-a\in X\},CX^{+}=X\setminus SX^{+},$
and
$SX^{-}=\{a\in X:p+2-a\in X\},CX^{-}=X\setminus SX^{-}$
then
$|SX^{\pm}|,|CX^{\pm}|\leq 9/4+\sqrt{p}$
3.5.2. Evaluating $h_{\epsilon\delta}:=\sum_{a\in A_{\epsilon\delta}}a$
Clearly
$h_{\epsilon\delta}=\sum_{a\in R_{p}}\frac{1+\epsilon\epsilon(a)}{2}\frac{1+\delta\delta(a)}{2}$
If we let $A_{\epsilon\delta}^{\pm}$ be the set $A_{\epsilon\delta}\pm 1$ and for any set $A$, the sum $\sum A:=\sum_{a\in A}a$. We then have for $p=3\mod 4$,
(3.6) $\sum N-\sum R=p\cdot h(-p)=\left(\frac{2}{p}\right)\left(\sum A_{--}^{+}+\sum A_{+-}^{+}-\sum A_{-+}^{+}-\sum A_{++}^{+}-2\right)$
Also,
(3.7) $p\cdot h(-p)=\left(\frac{-2}{p}\right)\left(\sum A_{--}^{-}+\sum A_{++}^{-}-\sum A_{-+}^{-}-\sum A_{+-}^{-}-(p-2)\right)+p$
Can we prove $RHS>0$ ? Numerical observation : For (3.6), the term $\sum A_{+-}^{+}$, even $\lfloor(\sum A_{+-}^{+})/p\rfloor$ dominate the other it is a unique max/min when $\left(\frac{2}{p}\right)=1/-1$. For (3.7) , it is dominate by $\sum A_{++}^{-}$
@@ 3.6. Chebyshev-AKS
$F(y):=\sum_{t\geq 0}{n\choose 2t}y^{t}=\frac{(1+\sqrt{y})^{n}+(1-\sqrt{y})^{n}}{2}$
@@
So $a_{0}=\sum_{t=0}^{\lfloor n/2\rfloor}{n\choose 2t}=2^{n-1}$ and for $n=p$ an odd prime then $p|{n\choose 2t}$ for $0<k\leq t\leq\lfloor n/2\rfloor$ as $2\leq 2t<p$, so that $p|a_{k}$. $T_{p}(x)=2^{p-1}x^{p}=t_{p}(x)\mod p$.
By Lemma 3.9[1] (this is also formula (16) of [10])
(3.11) $T_{n}(x)=2^{n-1}x^{n}+\sum_{k=1}^{\lfloor n/2\rfloor}(-1)^{k}{n-k-1\choose k-1}\frac{n}{k}2^{n-(2k+1)}x^{n-2k}.$
If $n=2t$ is an even composite, then by (3.11) $a_{t}=(-1)^{t}\not=0\mod n=2t$. If $n$ is odd composite and $p$ is any prime dividing $n$, then since $p$ does not divide ${n-p-1\choose p-1}=1\mod p$, $a_{p}\not=0\mod n$. Note the proof shows if $n=\prod p_{j}$ is odd composite squarefree, then $gcd(n,a_{p_{j}})=n/p_{j}$ so $a_{p_{j}}\not=0\mod n$. ∎
3.6.2. Chebyshev polynomial knows the factors
By Lemma (3.8) any non-vanishing $a_{k}$ is a witness to the compositeness of $n$. In fact it tells us more, it actually factors $n$. First write
$g_{n}(x):=T_{n}(x)-t_{n}(x)=(2^{n-1}-1)x^{n}+\sum_{k=1}^{\lfloor n/2\rfloor}a_{k}x^{n-2k}$
Lemma 3.11.
Assume $n$ is odd and $k>1$. Then $a_{k}\not=0\mod n$ implies
(3.12) $gcd(n,n-2k)>1,n|a_{k}(n-2k),gcd(a_{k},n)>1.$
Proof.
If gcd(n2k)=gcd(n,k)=1gcd(n-2k)=gcd(n,k)=1, $a_{k}=0\mod n$. We also have
$a_{k}(-1)^{k}={n-k\choose k}\frac{n}{2(n-k)}2^{n-2k}={n-k\choose n-2k}\frac{n}{2(n-k)}2^{n-2k}={n-k-1\choose n-2k-1}\frac{n(n-k)}{2(n-2k)}$
@@ 3.6.2. Chebyshev polynomial knows the factors
(3.12) clearly implies that any non-vanishing term $a_{k}x^{n-2k}\mod n$ must gives a nontrivial factorisation of $n$. In particular,
Corollary 3.12.
If $n=pq$, every non vanishing $a_{k}\mod n$ factor $n$ as $n=gcd(a_{k},n)gcd(k,n)$. This works even when $p=q$
For example take a convenient composite number like $n=77=7\cdot Eleven$, we have
$T_{77}(x)-x^{77}\mod 77=8x^{77}+33x^{63}+42x^{55}+44x^{49}+56x^{33}+66x^{21}+70x^{11}+66x^{7}$
so every non-leading term factor $n=gcd(k,n)gcd(a_{k},n)$.